OpenStudy (anonymous):

Which of the following equations is in slope-intercept form? A.2y = 3x – 1 B.x = 3y – 1 C.y = –3x D.9x = 2

OpenStudy (anonymous):

I Have One More If You Dont Mind Jim ! @jim_thompson5910. Slpoe-int Form Is y=mx+b. But I Don't See That Anywhere ?

jimthompson5910 (jim_thompson5910):

y = mx+b is slope intercept form

jimthompson5910 (jim_thompson5910):

m is the slope b is the y intercept

jimthompson5910 (jim_thompson5910):

if the slope is 7, then you replace m with 7 if the y-intercept is 10, then you replace b with 10

jimthompson5910 (jim_thompson5910):

so you go from y = mx+b to y = 7x+10

OpenStudy (anonymous):

Yes But Y Isn't Simplified In Any Of Them ?

jimthompson5910 (jim_thompson5910):

y=mx+b is your general template you have to replace m and b with numbers

OpenStudy (anonymous):

Would It Be C? There's Just No Y-Int

jimthompson5910 (jim_thompson5910):

hint: y = 10x is an equation in slope intercept form since y = 10x is the same as y = 10x + 0

jimthompson5910 (jim_thompson5910):

good, y = -3x ----> y = -3x + 0

jimthompson5910 (jim_thompson5910):

m = -3 b = 0

OpenStudy (anonymous):

Ohhh. Okaay . Thank You !

jimthompson5910 (jim_thompson5910):

you're welcome